Solve the system by addition. 5x – 3y = 13 4x – 3y = 11
Answer by praseena(37)

Solve the system by addition.

5x – 3y = 13 ----------(1)
4x – 3y = 11 ----------(2)

subtract second equation from first equation
===> 5x -3y - (4x - 3y ) = 13 - 11
===> 5x - 3y -4x + 3y = 3
==> 1x = 3
==> x = 3

thus we got the value x = 3
now substitute this value any one of the given equations.
here let us put x = 3 in the first equation, 5x -3y

==> 5* 3 - 3*y = 13
==> 15 -3y = 13
==> 15 -13 = 3y
==> 2 = 3y
==> y = 2/3

so the solution for this pair of equations is,
x = 3
y = 2/3